OpenStudy (anonymous):

How can I prove this for every 2x2 Matrix A?

OpenStudy (anonymous):

\[\det(A) = \frac{ 1 }{ 2 } \left[\begin{matrix}tr(A) & 1 \\ tr(A^{2}) & tr(A)\end{matrix}\right]\]